# Share Your Attention: Transformer Weight Sharing via Matrix-based Dictionary Learning

### v1

@@ Abstract
Large language models (LLMs) have revolutionized AI applications, yet their high computational and memory demands hinder their widespread deployment. Existing compression techniques focus on intra-block optimizations (e.g., low-rank approximation or attention head pruning), while the repetitive layered structure of transformers implies significant inter-block redundancy—a dimension largely unexplored beyond key-value (KV) caching. Inspired by dictionary learning in convolutional networks, we propose a framework for structured weight sharing across transformer layers. Our approach decomposes attention projection matrices (Q, K, V, O) into shared dictionary atoms, reducing the attention module’s parameters by 66.7% (e.g., 226.5M → 75M in a 700M-parameter model) while achieving on-par performance. Unlike complex methods requiring distillation or architectural changes, MASA (Matrix Atom Sharing in Attention) operates as a drop-in replacement—trained with standard optimizers—and represents each layer’s weights as linear combinations of shared matrix atoms. Experiments across scales (100M–700M parameters) show that MASA achieves better benchmark accuracy and perplexity than grouped-query attention (GQA), low-rank baselines and recently proposed Repeat-all-over/Sequential sharing at comparable parameter budgets. Ablation studies confirm robustness to the dictionary size and the efficacy of shared representations in capturing cross-layer statistical regularities. Extending to Vision Transformers (ViT), MASA matches performance metrics on image classification tasks with 66.7% fewer attention parameters. By combining dictionary learning strategies with transformer efficiency, MASA offers a scalable blueprint for parameter-efficient models without sacrificing performance. Finally, we investigate the possibility of employing MASA on large pretrained models to reduce their number of parameters without experiencing any significant drop in their performance.
@@ Introduction
Recently proposed methods like grouped-query attention (GQA) (Ainslie et al. 2023) and QK compression (e.g. LISA (Mu et al. 2024)) demonstrate the value of parameter reduction but operate within isolated layers or focus on particular projections inside attention module. Meanwhile, emerging approaches exploring cross-layer sharing—such as Repeat-all-over and Sequential parameter assignment strategies (Liu et al. 2024; Takase and Kiyono 2021)—reveal promising directions but suffer from performance degradation in complex reasoning tasks (Liao and Vargas 2024). Crucially, these methods lack a principled framework for capturing the statistical regularities across transformer layers.
Inspired by dictionary learning principles in convolutional networks (Mairal et al. 2009), we propose Matrix Atom Sharing in Attention (MASA), a novel framework that systematically exploits inter-block redundancy through structured weight sharing across transformer layers. Unlike prior sharing strategies that either enforce rigid weight tying or require complex distillation procedures, MASA decomposes attention projection matrices (Q, K, V, O) into shared dictionary atoms, enabling each layer’s weights to be represented as linear combinations of these atoms. This approach reduces attention module parameters by 66.7% (e.g., 226.5M → 75M in a 700M-parameter model) while maintaining competitive performance—achieving what previous parameter-sharing methods like Sequential-sharing and Repeat-all-over Sharing could not: consistent accuracy across diverse benchmarks and on-par (or better) performance than the original Transformer.
@@
    Parameter Efficiency with Performance Parity: MASA exceeds the performance of low-rank baselines, GQA, and recent Repeat-all-over/Sequential sharing approaches across language modeling (perplexity), reasoning, and knowledge benchmarks under the same (or higher) compression rate. Moreover, MASA with 66.7% less parameters in attention can match the performance of vanilla (uncompressed) Transformer for S, M, and L sizes.
@@
Figure 1: MASA framework: (Left) Independent dictionary pools for Q, K, V, O projections. (Middle) Per-block projection matrices synthesized via weighted combinations of shared dictionaries (example: Block l). All blocks share dictionary pools while using unique linear coefficients for each Transformer block.
Beyond language models, we demonstrate the broad applicability of MASA by extending it to Vision Transformers (ViTs), where it achieves strong performance on image classification tasks while compressing attention modules by 66.7%. Given the dominance of pretrained models in modern deployment pipelines, we further investigate MASA in training-free adaptation scenarios. Our experiments show that MASA incurs only marginal performance degradation upon parameter pruning, highlighting its robustness and practicality in resource-constrained settings. By unifying dictionary learning with architectural design in Transformers, MASA provides a principled and scalable framework for constructing parameter-efficient models without compromising accuracy. The rest of the paper details our proposed method and presents comprehensive evaluations across model scales (100M–700M parameters) on language and vision tasks. Finally, we conclude with applications to training-free adaptation, highlighting MASA’s potential for plug-and-play efficiency in pretrained ecosystems.
@@
To reduce inter-block redundancy, several works have explored reusing weight matrices. Weight tying between embedding and output layers is common (Press and Wolf 2017), and Universal Transformers (Dehghani et al. 2019) share parameters across time steps. More recently, MobileLLM (Liu et al. 2024) and Sequential-sharing (Takase and Kiyono 2021) apply deterministic patterns to share attention and FFN weights across layers.
@@
MASA extends this principle to Transformer weight matrices, treating each attention projection as a signal to be reconstructed from a shared dictionary. Unlike low-rank methods that impose global rank constraints, MASA allows for modular, projection-specific dictionaries and adaptive sparsity. The resulting decomposition is both expressive and highly parameter-efficient, achieving a higher compression ratio without performance loss.
@@
where in our case $\mathbf{W}=\begin{bmatrix}\operatorname*{vec}{\left(\mathbf{W}_{1}\right)}&\ldots&\operatorname*{vec}{\left(\mathbf{W}_{L}\right)}\end{bmatrix}\in\mathbb{R}^{d\cdot h\times L}$ is composed by stacking horizontally the vectorized versions of the model weights for the $L$ blocks in the network, $\mathbf{D}=\begin{bmatrix}\operatorname*{vec}{\left(\mathbf{D}_{1}\right)}&\ldots&\operatorname*{vec}{\left(\mathbf{D}_{S}\right)}\end{bmatrix}\in\mathbb{R}^{d\cdot h\times S}$ is the dictionary, where $\operatorname*{vec}{\left(\mathbf{D}_{s}\right)}\in\mathbb{R}^{d\cdot h},s=1,\ldots S$ represents the $s$-th matrix atom, $\mathbf{D}_{s}\in\mathbb{R}^{d\times h}$, in vectorized form and $S$ is the total number of dictionary atoms, while $\mathbf{C}\in\mathbb{R}^{S\times L}$ represents the linear coefficients of the representation.
@@
$\hat{\mathbf{W}}_{l}=\sum\limits_{s=1}^{S}c_{ls}\mathbf{D}_{s},\,\,\mbox{with}\,\,\mathbf{D}_{s}\in\mathbb{R}^{d\times h},c_{ls}\in\mathbb{R},$ (2)
where $c_{ls}$ is a scalar entry of the coefficient matrix $\mathbf{C}$. The above formulation describes our weight-sharing mechanism, where each individual weight matrix $\mathbf{W}_{l}$ is defined by a collection of shared weights (dictionary $\mathbf{D}$) and individual per block mixing coefficients ($\mathbf{c}_{l}=\begin{bmatrix}c_{l1}&\ldots&c_{lS}\end{bmatrix}\in\mathbb{R}^{S}$). By employing this strategy in the design of a transformer model, we can significantly reduce the network parameters, with the exact compression rate for a specific type of weight projection matrix computed as $r=1-\frac{S\left(d\cdot h+L\right)}{L\cdot d\cdot h}\approx 1-\frac{S}{L}$, with $S<L$ and $L<<d\cdot h$.
In dictionary learning, the optimal pair of the dictionary $\mathbf{D}$ and linear coefficients $\mathbf{C}$ are usually estimated by minimizing the approximation error
$\mathbf{D}^{\star},\mathbf{C}^{\star}=\operatorname*{arg\,min}_{\mathbf{D}\in\mathcal{D},\mathbf{C}\in\mathcal{C}}\left\|\mathbf{W}-\mathbf{D}\mathbf{C}\right\|_{F}^{2},$ (3)
@@
Similarly to MASA’s approach, as described in Eq. (2), given a set of pretrained weights $\left\{\mathbf{W}_{l}\right\}_{l=1}^{L}$ we aim to approximate each one as a linear combination of a collection of shared matrix components. However, unlike our training-from-scratch strategy, we don’t rely on the network loss and instead we aim for an analytical solution that minimizes the norm of the approximation error between the pretrained weights and the approximated ones. We further require that the shared matrix components are of unit-norm and orthogonal to each other, that is it holds $\operatorname*{tr}\left(\mathbf{D}_{i}^{\mathsf{T}}\mathbf{D}_{j}\right)=\delta_{ij}$ with $\delta$ the Kronecker delta function, while the linear coefficients are computed as $c_{ls}=\operatorname*{tr}\left(\mathbf{D}_{s}^{\mathsf{T}}\mathbf{W}_{l}\right)$. In other words, we are looking for a matrix basis of a subspace in $\mathbb{R}^{d\times h}$. The basis matrices, which constitute the basis, can be recovered as the minimizer of the following objective:
$\mathbf{D}_{1}^{\star},\ldots,\mathbf{D}_{S}^{\star}\!=\!\!\!\!\operatorname*{arg\,min}_{\begin{subarray}{c}\mathbf{D}_{s}\in\mathbb{R}^{d\times h}\\ \operatorname*{tr}\left(\mathbf{D}_{i}^{\mathsf{T}}\mathbf{D}_{j}\right)=\delta_{ij}\end{subarray}}\!\!\!\sum\limits_{l=1}^{L}\left\|\mathbf{W}_{l}-\sum\limits_{s=1}^{S}\operatorname*{tr}\left(\mathbf{D}_{s}^{\mathsf{T}}\mathbf{W}_{l}\right)\mathbf{D}_{s}\right\|_{F}^{2},$ (4)
where $\operatorname*{tr}\left(\cdot\right)$ denotes the matrix trace. Fortunately, the above minimization problem admits a closed-form solution (we refer to the appendix for a detailed derivation) that involves the eigenvectors corresponding to the $S$ largest eigenvalues of the matrix product $\mathbf{W}\mathbf{W}^{\mathsf{T}}$, with $\mathbf{W}$ defined as in Eq. (1).
@@
To enhance the fidelity of MASA in pretrained models without fine-tuning, we introduce a data-informed local refinement strategy that captures reconstruction residuals with compact, structured representations. After grouping blocks and computing shared dictionary atoms via Matrix PCA, we reconstruct each layer’s weights and compute the residual $\Delta\mathbf{W}_{l}=\mathbf{W}_{l}-\hat{\mathbf{W}}_{l}$. Instead of modeling $\Delta\mathbf{W}_{l}$ directly, we apply a Cholesky whitening transform based on calibration data, and approximate $\mathbf{L}_{l}\Delta\mathbf{W}_{l}$ with a low-rank matrix, where $\mathbf{L}_{l}$ is the Cholesky factor (Meyer 2023) of the input autocorrelation. This accounts for data geometry and improves approximation efficiency.
We further propose an adaptive rank allocation scheme that distributes the residual budget according to the role of each weight matrix in the attention computation graph. Motivated by the rank inequality $\text{rank}(\mathbf{A}\mathbf{B})\leq\min(\text{rank}(\mathbf{A}),\text{rank}(\mathbf{B}))$, we allocate more residual capacity to matrices with higher intrinsic rank (e.g., $\mathbf{W}_{q}$, $\mathbf{W}_{o}$) and less to those with structural constraints (e.g., $\mathbf{W}_{k}$, $\mathbf{W}_{v}$ in GQA/MQA). This asymmetric allocation ensures optimal use of the parameter budget under architectural imbalances. Detailed description of the proposed dynamic ranking algorithm is provided in the Appendix.
@@
- High compression: 66.7% reduction in attention parameters, achieved by employing $S=L/3$ shared matrices separately across each of Q, K, V, and O projections (denoted MASA-QKVO).
- Moderate compression: 50% reduction, where only Q, K, and V projections are defined using $S=L/3$ shared weights, while the O projections for each transformer block are left untouched (denoted MASA-QKV).
@@
Moreover, to enhance the stability and adaptability of the learned mixing factors $\mathbf{c}_{l}\in\mathbb{R}^{S}$ for each block $l$, we introduce a block-specific embedding-based parameterization. Specifically, each block is assigned a unique trainable embedding vector, which serves as input to a 3-layer MLP that predicts the corresponding coefficients $\mathbf{c}_{l}$. This over-parameterized formulation decouples the optimization dynamics of the mixing coefficients from direct, potentially unstable, updates, thereby reducing gradient fluctuations during training and promoting smoother convergence. Importantly, this design acts as an implicit regularization mechanism, guiding the model toward more stable configurations. After training, both MLP and embeddings are discarded, and only the final coefficient matrix C is retained for inference. This ensures no additional computational overhead at test time while preserving the benefits of smooth and more efficient training.
@@
We assess zero-shot performance across two benchmark families: multiple-choice reasoning and language modeling. We calculate average accuracy for PIQA (Bisk et al. 2019), HellaSwag (Zellers et al. 2019), MMLU (Hendrycks et al. 2020) and ARC Challenge (Clark et al. 2018) testsets. Also, we estimate perplexity for LAMBADA (Paperno et al. 2016) and WikiText (Merity et al. 2016). Detailed description for each benchmark can be found in the Appendix.

### v2

# Stamatios Lefkimmiatis1

@@ Abstract
Large language models (LLMs) have revolutionized AI applications, yet their high computational and memory demands hinder their widespread deployment. Existing compression techniques focus on intra-block optimizations (e.g., low-rank approximation or attention head pruning), while the repetitive layered structure of transformers implies significant inter-block redundancy—a dimension largely unexplored beyond key-value (KV) caching. Inspired by dictionary learning in convolutional networks, we propose a framework for structured weight sharing across transformer layers. Our approach decomposes attention projection matrices (Q, K, V, O) into shared dictionary atoms, reducing the attention module’s parameters by 66.7% (e.g., 226.5M → 75M in a 700M-parameter model) while achieving on-par performance. Unlike complex methods requiring distillation or architectural changes, MASA (Matrix Atom Sharing in Attention) operates as a drop-in replacement—trained with standard optimizers—and represents each layer’s weights as linear combinations of shared matrix atoms. Experiments across scales (100M–700M parameters) show that MASA achieves better benchmark accuracy and perplexity than grouped-query attention (GQA), low-rank baselines and recently proposed Repeat-all-over/Sequential sharing at comparable parameter budgets. Ablation studies confirm robustness to the dictionary size and the efficacy of shared representations in capturing cross-layer statistical regularities. Extending to Vision Transformers (ViT), MASA matches performance metrics on image classification tasks with 66.7% fewer attention parameters. By combining dictionary learning strategies with transformer efficiency, MASA offers a scalable blueprint for parameter-efficient models without sacrificing performance. Finally, we investigate the possibility of employing MASA on large pretrained models to reduce their number of parameters without experiencing any significant drop in their performance. Code will be available at https://github.com/mts-ai/MASA
**footnotetext: This work was accepted and presented at AAAI 2026
@@ Introduction
Recently proposed methods like grouped-query attention (GQA) (Ainslie et al. 2023) and QK compression (e.g. LISA (Mu et al. 2024)) demonstrate the value of parameter reduction but operate within isolated layers or focus on particular projections inside attention module. Meanwhile, emerging approaches exploring cross-layer sharing—such as Repeat-all-over (Liu et al. 2024) and Sequential parameter (Takase and Kiyono 2021) assignment strategies—reveal promising directions but suffer from performance degradation in reasoning tasks (Liao and Vargas 2024). Crucially, these methods lack a principled framework for capturing the statistical regularities across transformer layers.
Inspired by dictionary learning principles in convolutional networks (Mairal et al. 2009), we propose Matrix Atom Sharing in Attention (MASA), a novel framework that systematically exploits inter-block redundancy through structured weight sharing across transformer layers. Unlike prior sharing strategies that either enforce rigid weight tying or require complex distillation procedures, MASA decomposes attention projection matrices (Q, K, V, O) into shared dictionary atoms, enabling each layer’s weights to be represented as linear combinations of these atoms. This approach reduces attention module parameters by 66.7% (e.g., 226.5M → 75M in a 700M-parameter model) while maintaining competitive performance—achieving what previous parameter-sharing methods like Sequential-sharing (Takase and Kiyono 2021) and Repeat-all-over Sharing could not: consistent accuracy across diverse benchmarks and on-par (or better) performance than the original Transformer.
@@
    Parameter Efficiency with Performance Parity: MASA exceeds the performance of low-rank baselines, GQA, and recent Repeat-all-over/Sequential sharing approaches across language modeling (perplexity), reasoning, and knowledge benchmarks under the same (or higher) compression rate. Moreover, MASA with 66.7% less parameters in attention can match the performance of vanilla (uncompressed) Transformer for S, M, L sizes.
@@
Figure 1: MASA framework: (Left) Independent dictionary pools for Q, K, V, O projections. (Middle) Per-block projection matrices synthesized via weighted combinations of shared dictionaries. All blocks share dictionary pools while using unique linear coefficients for each Transformer block.
Beyond language models, we demonstrate the broad applicability of MASA by extending it to Vision Transformers (ViTs), where it achieves strong performance on image classification tasks while compressing attention modules by 66.7%. Given the dominance of pretrained models in modern deployment pipelines, we further investigate MASA in training-free adaptation scenarios. Our experiments show that MASA incurs only marginal performance degradation upon parameter pruning, highlighting its robustness and practicality in resource-constrained settings. By unifying dictionary learning with architectural design in Transformers, MASA provides a principled and scalable framework for constructing parameter-efficient models without compromising accuracy. The rest of the paper details our proposed method and presents comprehensive evaluations across model scales (100M–700M parameters) on language and vision tasks. Finally, we conclude with applications to training-free adaptation, highlighting MASA’s potential for plug-and-play efficiency in pretrained models.
@@
To reduce inter-block redundancy, several works have explored reusing weight matrices. Weight tying between embedding and output layers is common (Press and Wolf 2017), and Universal Transformers (Dehghani et al. 2019) share parameters across time steps. More recently, Repeat-all-over (Liu et al. 2024) and Sequential-sharing apply deterministic patterns to share attention and FFN weights across layers.
@@
MASA extends this principle to Transformer weight matrices, treating each attention projection as a signal to be reconstructed from a shared dictionary. Unlike low-rank methods that enforce a uniform rank constraint across all layers and projection types, MASA learns separate dictionaries of shared matrix atoms for each attention projection and represents each layer’s weights as adaptive linear combinations of these atoms, allowing the effective rank to vary by layer and projection type. This flexible decomposition achieves higher compression ratios without sacrificing performance.
@@
where in our case $\mathbf{W}=\begin{bmatrix}\operatorname{vec}(\mathbf{W}_{1})&\cdots&\operatorname{vec}(\mathbf{W}_{L})\end{bmatrix}\in\mathbb{R}^{d\cdot h\times L}$ is composed by stacking horizontally the vectorized versions of the model weights for the $L$ blocks in the network, $\mathbf{D}=\begin{bmatrix}\operatorname{vec}(\mathbf{D}_{1})&\cdots&\operatorname{vec}(\mathbf{D}_{S})\end{bmatrix}\in\mathbb{R}^{d\cdot h\times S}$ is the dictionary, where $\operatorname{vec}(\mathbf{D}_{s})\in\mathbb{R}^{d\cdot h}$, $s=1,\dots,S$ represents the $s$-th matrix atom, $\mathbf{D}_{s}\in\mathbb{R}^{d\times h}$, in vectorized form and $S$ is the total number of dictionary atoms, while $\mathbf{C}\in\mathbb{R}^{S\times L}$ represents the linear coefficients of the representation.
@@
$\hat{\mathbf{W}}_{l}=\sum_{s=1}^{S}c_{ls}\mathbf{D}_{s},\text{ with }\mathbf{D}_{s}\in\mathbb{R}^{d\times h},c_{ls}\in\mathbb{R},$ (2)
where $c_{ls}$ is a scalar entry of the coefficient matrix $\mathbf{C}$. The above formulation describes our weight-sharing mechanism, where each individual weight matrix $\mathbf{W}_{l}$ is defined by a collection of shared weights (dictionary $\mathbf{D}$) and individual per block mixing coefficients ($\mathbf{c}_{l}=[c_{l1}\;\cdots\;c_{lS}]^{\top}\in\mathbb{R}^{S}$). By employing this strategy in a transformer model, we can significantly reduce the network parameters, with the exact compression rate for a specific type of weight projection matrix computed as $r=1-\frac{S(d\cdot h+L)}{L\cdot d\cdot h}\approx 1-\frac{S}{L}$, with $S<L$ and $L<d\cdot h$.
In dictionary learning, the optimal pair of the dictionary $\mathbf{D}$ and linear coefficients $\mathbf{C}$ are usually estimated by minimizing the approximation error.
$\mathbf{D}^{*},\mathbf{C}^{*}=\mathop{\mathrm{arg\,min}}_{\mathbf{D}\in\mathcal{D},\,\mathbf{C}\in\mathcal{C}}\left\|\mathbf{W}-\mathbf{D}\mathbf{C}\right\|_{F}^{2},$ (3)
@@
Similarly to MASA’s approach, as described in Eq. (2), given a set of pretrained weights $\{\mathbf{W}_{l}\}_{l=1}^{L}$ we aim to approximate each one as a linear combination of a collection of shared matrix components. However, unlike our training-from-scratch strategy, we don’t rely on the network loss and instead we aim for an analytical solution that minimizes the norm of the approximation error between the pretrained weights and the approximated ones. We further require that the shared matrix components are of unit-norm and orthogonal to each other, so it holds $\operatorname{tr}\left(\mathbf{D}_{i}^{\top}\mathbf{D}_{j}\right)=\delta_{ij}$ with $\delta$ the Kronecker delta function, while the linear coefficients are computed as $c_{ls}=\operatorname{tr}\left(\mathbf{D}_{s}^{\top}\mathbf{W}_{l}\right)$. Thus, we are looking for a matrix basis of a subspace in $\mathbb{R}^{d\times h}$. The basis matrices, which constitute the basis, can be recovered as the minimizer of the following objective:
$\centering\mathbf{D}_{1}^{*},..,\mathbf{D}_{S}^{*}=\mathop{\mathrm{arg\,min}}_{\begin{subarray}{c}\mathbf{D}_{s}\in\mathbb{R}^{d\times h}\\ \operatorname{tr}\left(\mathbf{D}_{i}^{\top}\mathbf{D}_{j}\right)=\delta_{ij}\end{subarray}}\sum_{l=1}^{L}\left\|\mathbf{W}_{l}-\sum_{s=1}^{S}\operatorname{tr}\left(\mathbf{D}_{s}^{\top}\mathbf{W}_{l}\right)\mathbf{D}_{s}\right\|_{F}^{2},\@add@centering$ (4)
where $\operatorname{tr}(\cdot)$ denotes the matrix trace. Fortunately, the above minimization problem admits a closed-form solution (we refer to the appendix for a detailed derivation) that involves the eigenvectors corresponding to the $S$ largest eigenvalues of the matrix product $\mathbf{W}\mathbf{W}^{\top}$, with $\mathbf{W}$ defined as in Eq. (1).
@@
To enhance the fidelity of MASA in pretrained models without fine-tuning, we introduce a data-informed local refinement strategy that captures reconstruction residuals with compact, structured representations. After grouping blocks and computing shared dictionary atoms via Matrix PCA, we reconstruct each layer’s weights and compute the residual $\Delta\mathbf{W}_{l}=\mathbf{W}_{l}-\hat{\mathbf{W}}_{l}$. Instead of modeling $\Delta\mathbf{W}_{l}$ directly, we apply a Cholesky whitening transform based on calibration data, and approximate $L_{l}\Delta\mathbf{W}_{l}$ with a low-rank matrix, where $L_{l}$ is the Cholesky factor (Meyer 2023) of the input autocorrelation. This accounts for data geometry and improves approximation efficiency.
We further propose an adaptive rank allocation scheme that distributes the residual budget according to the role of each weight matrix in the attention computation. Motivated by the rank inequality $\operatorname{rank}(\mathbf{A}\mathbf{B})\leq\min\{\operatorname{rank}(\mathbf{A}),\operatorname{rank}(\mathbf{B})\}$, we allocate more residual capacity to matrices with higher intrinsic rank (e.g., $\mathbf{W}_{q}$, $\mathbf{W}_{o}$) and less to those with structural constraints (e.g., $\mathbf{W}_{k}$, $\mathbf{W}_{v}$ in GQA/MQA). This asymmetric allocation ensures optimal use of the parameter budget under architectural imbalances. Detailed description of the proposed dynamic ranking algorithm is provided in the Appendix.
@@
1. High compression: 66.7% reduction in attention parameters, achieved by employing $S=L/3$ shared matrices separately across each of Q, K, V, and O projections (denoted MASA-QKVO).
2. Moderate compression: 50% reduction, where only Q, K, and V projections are defined using $S=L/3$ shared weights, while the O projections for each transformer block are left untouched (denoted MASA-QKV).
@@
Moreover, to enhance the stability and adaptability of the learned mixing factors $\mathbf{c}_{l}\in\mathbb{R}^{S}$ for each block $l$, we introduce a block-specific embedding-based parameterization. Each block is assigned a unique trainable embedding vector, which serves as input to a 3-layer MLP that predicts the corresponding coefficients $\mathbf{c}_{l}$. This parameterized formulation decouples the optimization dynamics of the mixing coefficients from direct, potentially unstable, updates, thereby reducing gradient fluctuations during training and promoting smoother convergence. Importantly, this design acts as an implicit regularization mechanism, guiding the model toward more stable configurations. After training, both MLP and embeddings are discarded, and only the final coefficient matrix $\mathbf{C}$ is retained for inference. This ensures no additional computational overhead at test time while preserving the benefits of smooth and more efficient training.
@@
We assess zero-shot performance across two benchmark families: multiple-choice reasoning and language modeling. We compute the average accuracy across 8 test sets, including PIQA (Bisk et al. 2019), HellaSwag (Zellers et al. 2019), and ARC Challenge (Clark et al. 2018). Also, we estimate perplexity for LAMBADA (Paperno et al. 2016) and WikiText (Merity et al. 2016). Detailed description for each benchmark can be found in the Appendix.